# Mean uniformly stable function and its application to almost sure stability analysis of randomly switched time-varying systems

### v2

@@ Abstract
This paper investigates uniform almost sure stability of randomly switched time-varying systems. Mode-dependent indefinite multiple Lyapunov functions (iMLFs) are introduced to assess stability properties of diverse time-varying subsystems. To realize the stability conditions establishment based on iMLFs, we present a novel condition so-called mean uniformly stable function for time-varying parameters of iMLFs’ derivatives. Our approach provides a random perspective, which makes iMLFs suits for random switched time-varying systems. Moreover, the MUSF condition revealed a fact that each time-varying subsystem staying mean bounded during its corresponding sojourn time interval is a precondition for whole system almost sure stable. The combination of iMLFs and MUSFs, to some extent, preforms a flexibility to accommodate stability analysis with unstable subsystems and stable but no-exponentially decay subystems. Numerical examples are provided to demonstrate the effectiveness and advantages of our approach.
@@ I Introduction
Stability plays a crucial and fundamental role in the dynamic analysis and control design of systems. The main concepts of stability for randomly switched systems include: stability in probability, $p$th moment stability and almost sure stability [1, 4, 5]. Stability in probability characterizes the convergence behaviors of states through a probability measure, requiring that the events of stable state trajectories occur with a nonzero probability. In contrast, the $p$th moment stability emphasizes the average behavior of trajectories, representing stable states through mathematical expectation. This stability concept simplifies stochastic analysis problems into deterministic ones, making it widely applicable in the analysis and synthesis of stochastic systems [6, 7, 8]. Notably, $p$th moment stability can imply the stability in probability, some authors even express stability in probability via integral of the mean square ($2$th moment) of the states instead of probability measure (see Definition 2.1 of [9] and Definition 1 of [10]). In addition, almost sure stability requires that the events of the stable state trajectory occur with probability 1. As pointed out by [9], almost sure stability is of significant practical importance, as it is the sample paths—rather than the $p$th moments of the state process—that are observed in practice.
Given its practical utility in various stochastic systems, almost sure stability serves as the primary focus of investigation in this work. For the linear Markov switched time-invariant systems, the equivalence of three types of stochastic stability has been proved in [9]. However, for the nonlinear randomly switched systems, such equivalence no longer holds [4, 11]. In this work, we aim to establish sufficient condition for uniform almost sure stability in randomly switched systems, considering all the subsystems are time-varying. In doing so, we will explore what stochastic properties are satisfied by the subsystems to ensure the almost sure stability of the whole system.
@@
where $\dot{V}_{i}(t,x)=\frac{\partial V_{i}(t,x(t))}{\partial t}+\frac{\partial V_{i}(t,x(t))}{\partial x}$. The prototype of this form can be traced back to the research of time-varying systems in Krasovskii’s monograph [12] in 1959. In Theorem 10.3 of [12], Krasovskii explored the idea of replacing the negative defined constant parameter of the time derivative of Lyapunov function with a time-varying function, i.e.,
@@
For randomly switched systems, its outstanding advantage is describing different energy decay rates of subsystems, even some of them are unstable [22, 23, 24]. For switched time-varying systems, it’s natural to combine MLFs (3) and iLFs (2) to investigate stability [25, 26, 27, 28, 29, 30]. However, in the literature mentioned above iMLFs method assumes derivatives of Lyapunov functions share a common time-varying parameter, that is,
@@
  • This novel constraint requires integrals of $\lambda_{i}(t)$ between every two successive switching points are mean bounded. Our evaluation actually is based on the expected length of the interval confined instead of its real length, which no need to know switching time points $t_{k}$ and $t_{k+1}$.
@@ II Model formulation and preliminaries
where $x(t)\in\mathbb{R}^{n}$ is state, $r(t)\in\mathbb{S}\subset\mathbb{N}^{+}$ is the switching signal and also a cádlág stochastic process, where $\mathbb{S}$ is set of the switch modes. For all $r(t)=i\in\mathbb{S}$, $f_{i}$ satisfy the locally Lipschitz condition and $f_{i}(t,0,0)=0$. For simplify, we note $r(t_{0})=r_{0}$. The set of switching moments is defined by $\mathcal{T}=\{t_{0}\leq t_{1}\leq t_{2}\leq\cdots\leq t_{k}\leq\cdots,\ k\in\mathbb{N}^{+}\}$. $N_{i}(t,s)$ represents the occurrence counts of $i$th subsystem during $[s,t]$, and $N_{ij}(t,s)$ is the counts of transitions from $i$th to $j$th subsystem and. The counting function
$N(t,s)=\sum_{i\in\mathbb{S}}N_{i}(t,s)=\sum_{i,j\in\mathbb{S}}N_{ij}(t,s)$ (9)
is the total switch counts. For $k\in\mathbb{N}^{+},\ i,j\in\mathbb{S}$, $S(k)=t_{k}-t_{k-1}$, $S_{i}(k)$ and $S_{ij}(k)$ are $k$th sojourn time, the $k$th sojourn time of mode $i$, and the sojourn time of mode $j$ at the $k$th transition from mode $i$ to mode $j$, respectively. Over the time interval $[s,t]$, $T_{i}(t,s)=\sum_{k=1}^{N_{i}(t,s)}S_{i}(k)$ and $T_{ij}(t,s)=\sum_{k=1}^{N_{ij}(t,s)}S_{ij}(k)$ are the total sojourn time of mode $i$ and the total sojourn time from mode $i$ to mode $j$, respectively.
@@
    $\mathbb{P}\left[r_{k}=i,S(k)\leq t|\cup_{l=1}^{k-1}\left\{r_{l},S(l)\right\}\right]=\mathbb{P}\left[r_{k}\!=\!i,S(k)\!\leq\!t|r_{k-1}\right],i\in\mathbb{S},t\geq t_{0}.$
@@
    $\pi_{i}=\frac{\bar{\pi}_{i}m_{i}}{\sum_{j\in\mathbb{S}}\bar{\pi}_{j}m_{j}},\ i\in\mathbb{S},$
@@ III Main Results
In this section, we will present a novel approach called mean uniformly functions (MUSFs) condition for iMLFs, aimed at establishing sufficient conditions for almost sure GUAS and GES. This section will demonstrate the importance of MUSFs in the stability analysis of randomly switching time-varying systems.
@@ III-A An illustrative example
To verify its stability, we introduce the Lyapunov exponent
$\lim_{t\rightarrow\infty}\frac{1}{t}\ln|x(t)|$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln|x_{0}|+\lim_{t\rightarrow\infty}\frac{1}{t}\int_{0}^{t}a_{r(s)}(s)\mathrm{d}s$
@@
Let’s consider that all the functions $a_{i}(s)$ are the constants, that is $a_{i}(s)=\bar{a}_{i}\in\mathbb{R}$ for all $i\in\mathbb{S}$, then $\int_{t_{i_{k}}}^{t_{i_{k+1}}}a_{i}(s)\mathrm{d}s=\bar{a}_{i}(t_{i_{k+1}}-t_{i_{k}})=\bar{a}_{i}S_{i}(k)$. By the strong law of large number [34] and ergodicity of $r$, we have
$\lim_{t\rightarrow\infty}\sum_{i\in\mathbb{S}}\frac{N_{i}(t,0)}{t}\frac{1}{N_{i}(t,0)}\sum_{k=0}^{N_{i}(t,0)}a_{i}S_{i}(k)=\sum_{i\in\mathbb{S}}\frac{\pi_{i}}{m_{i}}a_{i}m_{i}=\sum_{i\in\mathbb{S}}\pi_{i}a_{i},\ \mathrm{a.s.}$ (12)
If $\sum_{i\in\mathbb{S}}\pi_{i}a_{i}<0$, system will be GES a.s. In the above example, irreducibility of the semi-Markov process does play a critical in deriving almost sure stability. Meanwhile, we can notice that for each mode $i$ sequence $\left\{\bar{a}_{i}S_{i}(k)\right\}_{k\in\mathbb{N}^{+}}$ is an independent, identically distributed (i.i.d.) process and also a wide-sense stationary process. This fact also plays a crucial role in deducing (12). For each $i$, we define a random variable for sequence $\{\bar{a}_{i}S_{i}(k)\}$ which is its time average:
@@
We then use Chebychev’s inequality to verify (12) holds:
@@
$\mathrm{Var}(A_{S}^{i})$ $=\mathbb{E}\left[(A_{S}^{i}-\bar{A}_{S}^{i})^{2}\right]$
@@
$\lim_{t\rightarrow\infty}\frac{1}{t}\ln|x(t)|$ $=\lim_{t\rightarrow\infty}\sum_{i\in\mathbb{S}}\frac{N_{i}(t,0)}{t}\frac{1}{N_{i}(t,0)}\sum_{k=0}^{N_{i}(t,0)}\int_{t_{i_{k}}}^{t_{i_{k+1}}}a_{i}(s)\mathrm{d}s,$
@@
Define a random variable $A_{y}^{i}$ from $\{Y_{k}^{i}\}_{k\in\mathbb{N}^{+}}$ as
$A_{y}^{i}$ $=\lim_{N_{i}\rightarrow\infty}\frac{1}{N_{i}}\sum_{k=0}^{N_{i}}Y_{k}^{i},$
$\bar{A}_{y}^{i}$ $=\mathbb{E}[A_{y}^{i}]=\mathbb{E}\left[\lim_{N_{i}\rightarrow\infty}\frac{1}{N_{i}}\sum_{k=0}^{N_{i}}Y_{k}^{i}\right]$
$=\lim_{N_{i}\rightarrow\infty}\frac{1}{N_{i}}\sum_{k=0}^{N_{i}}\mathbb{E}[Y_{k}^{i}]$
$=\mathbb{E}[Y_{k}^{i}]=M_{i}$
where $N_{i}=N_{i}(t,t_{0})$, $N_{i}\rightarrow\infty$ as $t\rightarrow\infty$. We claim that the sequence $\{Y^{i}_{k}\}_{k\in\mathbb{N}^{+}}$ is ergodic, that is
$A^{i}_{y}=\lim_{t\rightarrow\infty}\frac{1}{N_{i}}\sum_{k=0}^{N_{i}}b_{i}(S_{i}(k))=\mathbb{E}\left[b_{i}(S_{i}(k))\right]=M_{i}=\bar{A}_{y}^{i},\ \mathrm{a.s.}$ (15)
[…]
$\mathbb{P}\left[|A_{y}^{i}-\bar{A}_{y}^{i}|<\varepsilon\right]\geq 1-\frac{\mathrm{Var}(A_{y}^{i})}{\varepsilon^{2}}$ (16)
where $\mathrm{Var}(A_{y}^{i})$ is variance of $A_{y}^{i}$. For our random process to be ergodic, we need the probability in (16) to be 1 no matter the value of $\varepsilon$ is. For this purpose, we need the $\mathrm{Var}(A_{y}^{i})=0$. Now, we calculate the variance:
$\mathrm{Var}(A_{y}^{i})$ $=\mathbb{E}\left[(A_{y}^{i}-\bar{A}_{y}^{i})^{2}\right]$
$=\mathbb{E}\left[\lim_{N_{i}\rightarrow\infty}(\frac{1}{N_{i}}\sum_{k=0}^{N_{i}}Y_{k}^{i}-M_{i})(\frac{1}{N_{i}}\sum_{l=0}^{N_{i}}Y_{l}^{i}-M_{i})\right]$
$=\lim_{N_{i}\rightarrow\infty}\left(\frac{1}{N_{i}}\right)^{2}\sum_{k=0}^{N_{i}}\sum_{l=0}^{N_{i}}\mathbb{E}\left[(Y_{k}^{i}-M_{i})(Y_{l}^{i}-M_{i})\right]$
$=\lim_{N_{i}\rightarrow\infty}\left(\frac{1}{N_{i}}\right)^{2}\sum_{k=0}^{N_{i}}\sum_{l=0}^{N_{i}}\left(\mathbb{E}\left[Y_{k}^{i}Y_{l}^{i}\right]-M_{i}^{2}\right)$
$=\lim_{N_{i}\rightarrow\infty}\left(\frac{1}{N_{i}}\right)^{2}\sum_{k=0}^{N_{i}}\sum_{l=0}^{N_{i}}\left(\mathbb{E}\left[Y_{k}^{i}\right]\mathbb{E}\left[Y_{l}^{i}\right]-M_{i}^{2}\right)$
$=0,$
the 4th equation follows from the independence property of $\{Y_{k}^{i}\}_{k\in\mathbb{N}^{+}}$ and the 5th equation follows from $\mathbb{E}[Y^{i}_{k}]=M_{i}$. Then, (16) holds with probability 1, the sequence $\{Y_{k}^{i}\}_{k\in\mathbb{N}^{+}}$ is ergodic. Therefore, the Lyapunov exponent (11) follows from (15) that
$\lim_{t\rightarrow\infty}\frac{1}{t}\ln|x(t)|=$ $\lim_{t\rightarrow\infty}\sum_{i\in\mathbb{S}}\frac{N_{i}(t,0)}{t}\frac{1}{N_{i}(t,0)}\sum_{k=0}^{N_{i}(t,0)}\int_{t_{i_{k}}}^{t_{i_{k+1}}}a_{i}(s)\mathrm{d}s$
$\leq$ $\lim_{t\rightarrow\infty}\sum_{i\in\mathbb{S}}\frac{N_{i}(t,0)}{t}\frac{1}{N_{i}(t,0)}\sum_{k=0}^{N_{i}(t,0)}b_{i}(S_{i}(k))$
@@
When $\sum_{i\in\mathbb{S}}\frac{\pi_{i}}{m_{i}}M_{i}<0$, system is GES a.s. By this mean, we derive a sufficient conditions of GES a.s. for time-varying systems.
@@ III-B Indefinite multiple Lyapunov functions and mean uniformly functions
We aim to use iMLFs and properties of randomly switching signals to get sufficient conditions of GUAS a.s. and GEAS a.s. Thus, we introduce following iMLFs scheme.
@@
    where $\dot{V}_{i}(t,x)=\frac{\partial V_{i}(t,x(t))}{\partial t}+\frac{\partial V_{i}(t,x(t))}{\partial x}$;
@@
For each mode $i$, function $\lambda_{i}(t)\in\mathcal{PC}(\mathbb{R}^{+};\mathbb{R})$ is said to be a mean uniformly stable function (MUSF), if for corresponding random variable $S_{i}>0$ with $\mathbb{E}[S_{i}]<\infty$ there exists a continuous function $\varphi_{i}(s)\in\mathcal{C}(\mathbb{R}^{+};\mathbb{R})$ such that $\mathbb{E}\left[\varphi_{i}(S_{i})\right]<\infty$ and
$\int_{t}^{t+S_{i}}\lambda_{i}(h)\mathrm{d}h\leq\varphi_{i}(S_{i}),\ \forall t\geq t_{0}.$ (18)
@@
We name ’mean uniformly stable functions’ in Definition 3 for three main reasons. Firstly, each integral of $\lambda_{i}(t)$ is estimated only and exclusively during the relative subsystem’s sojourn time, which is a random variable. To capture these cumulative variations of $\lambda_{i}(t)$, we employ the expectation to quantify its integral, introducing a condition that operates in a random sense. Secondly, due to the piecewise continuity of function $\lambda_{i}(t)$, its integral is continuous, and existence of a continuous function $\varphi_{i}$ that bounds this integral is reasonable. As shown in (18), function $\varphi_{i}$ is independent on time but dependent on sojourn time intervals, which performs the uniformity in time. Third, The Definitions 3 implies that all subsystems are mean bounded during their respective sojourn time intervals. Each time-varying parameter $\lambda_{i}(t)$ performs mean boundedness during its sojourn time. Although functions $\lambda_{i}(t)$ are allowed to take positive value, MUSFs can still ensure the boundedness in mean sense of their integrals.

### v3

@@ Abstract
This paper investigates uniform almost sure stability of randomly switched time-varying systems. Mode-dependent indefinite multiple Lyapunov functions (iMLFs) are introduced to assess stability properties of diverse time-varying subsystems. To realize the stability conditions establishment based on iMLFs, we present a novel condition so-called mean uniformly stable function for time-varying parameters of iMLFs’ derivatives. Our approach provides a probabilistic perspective, making iMLFs well-suited for randomly switched time-varying systems. Moreover, the MUSF condition reveals an essential insight: ensuring that each time-varying subsystem remains mean-bounded during its corresponding sojourn time interval is a prerequisite for the almost sure stability of the entire system. Additionally, the combination of iMLFs and MUSFs is able to accommodate stability analysis scenarios where some subsystems are unstable or exhibit non-exponential decay. Numerical examples are provided to demonstrate the effectiveness and advantages of our approach
@@
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. DOI: 10.1109/TAC.2025.3580631
@@ I Introduction
Stability plays a crucial and fundamental role in the dynamic analysis and control design of systems. The main concepts of stability for randomly switched systems include: stability in probability, $p$th moment stability and almost sure stability [1, 4, 5]. Stability in probability characterizes the convergence behaviors of states through a probability measure, requiring that the events of stable state trajectories occur with a nonzero probability. In contrast, the $p$th moment stability emphasizes the average behavior of trajectories, representing stable states through mathematical expectation. This stability concept simplifies stochastic analysis problems into deterministic ones, making it widely applicable in the analysis and synthesis of stochastic systems [6, 7, 8]. Notably, $p$th moment stability can imply the stability in probability, some authors even express stability in probability via integral of the mean square ($2$th moment) of the states instead of probability measure (see Definition 2.1 of [9] and Definition 1 of [10]). Beyond these two concepts, almost sure stability offers a more practical perspective: it requires that the events of the stable state trajectory occur with probability 1. As pointed out by [9], almost sure stability is of significant practical importance, as it is the sample paths rather than the $p$th moments of the state process that are observed in practice.
Given its practical utility in various stochastic systems, almost sure stability serves as the primary focus of investigation in this work. For the linear Markov switched time-invariant systems, [9] has confirmed that $p$th moment stability implies almost sure stability but the reverse is not true. For the nonlinear randomly switched systems, there is no obvious inclusion relationship between these two stability concepts [4, 11]. In this work, we aim to investigate uniformly almost surely stable property of nonlinear randomly switched systems, considering all the subsystems are time-varying. In doing so, we will explore what stochastic properties are satisfied by the subsystems to ensure the almost sure stability of the whole system.
@@
where $\dot{V}_{i}(t,x)=\frac{\partial V_{i}(t,x)}{\partial t}+\frac{\partial V_{i}(t,x)}{\partial x}f_{i}(t,x)$. The prototype of this form can be traced back to the research of time-varying systems in Krasovskii’s monograph [12] in 1959. In Theorem 10.3 of [12], Krasovskii explored the idea of replacing the negative defined constant parameter of the time derivative of Lyapunov function with a time-varying function, i.e.,
@@
For randomly switched systems, its outstanding advantage is describing different energy decay rates of subsystems, even when some of them are unstable [22, 23, 24]. For switched time-varying systems, it’s natural to combine MLFs (3) and iLFs (2) to investigate stability [25, 26, 27, 28, 29, 30]. However, in the literature mentioned above iMLFs method assumes derivatives of Lyapunov functions share a common time-varying parameter, that is,
@@
  • This novel constraint requires integrals of $\lambda_{i}(t)$ between every two successive switching points are mean bounded. Our evaluation actually is based on the expected length of the interval confined instead of its real length, without the need to know switching time points $t_{k}$ and $t_{k+1}$.
@@ II Model formulation and preliminaries
where $x(t)\in\mathbb{R}^{n}$ is state, $r(t)\in\mathbb{S}\subset\mathbb{N}^{+}$ is the switching signal and also a càdlàg stochastic process, where $\mathbb{S}$ is set of the switch modes. For all $r(t)=i\in\mathbb{S}$, $f_{i}$ satisfy the locally Lipschitz condition and $f_{i}(t,0,0)=0$. For simplify, we note $r(t_{0})=r_{0}$. The set of switching moments is defined by $\mathcal{T}=\{t_{0}\leq t_{1}\leq t_{2}\leq\cdots\leq t_{k}\leq\cdots,\ k\in\mathbb{N}^{+}\}$. $N_{i}(t,s)$ represents the occurrence counts of $i$th subsystem during $[s,t]$, and $N_{ij}(t,s)$ is the counts of transitions from $i$th to $j$th subsystem and. The counting function $N(t,s)=\sum_{i\in\mathbb{S}}N_{i}(t,s)=\sum_{i,j\in\mathbb{S}}N_{ij}(t,s)$ is the total switch counts. For $k\in\mathbb{N}^{+},\ i,j\in\mathbb{S}$, $S(k)=t_{k}-t_{k-1}$, $S_{i}(k)$ and $S_{ij}(k)$ are $k$th sojourn time, the $k$th sojourn time of mode $i$, and the sojourn time of mode $j$ at the $k$th transition from mode $i$ to mode $j$, respectively. Over the time interval $[s,t]$, $T_{i}(t,s)=\sum_{k=1}^{N_{i}(t,s)}S_{i}(k)$ and $T_{ij}(t,s)=\sum_{k=1}^{N_{ij}(t,s)}S_{ij}(k)$ are the total sojourn time of mode $i$ and the total sojourn time from mode $i$ to mode $j$, respectively.
@@
    $\mathbb{P}\left[r_{k}=i,S(k)\leq t|\cup_{l=1}^{k-1}\left\{r_{l},S(l)\right\}\right]$
    $=\mathbb{P}\left[r_{k}\!=\!i,S(k)\!\leq\!t|r_{k-1}\right],i\in\mathbb{S},t\geq t_{0}.$
@@
    $\vspace{-0.2cm}\pi_{i}=\frac{\bar{\pi}_{i}m_{i}}{\sum_{j\in\mathbb{S}}\bar{\pi}_{j}m_{j}},\ i\in\mathbb{S},$
@@ III Main Results
In this section, we will present a novel approach called mean uniformly functions (MUSFs) condition for iMLFs to establish sufficient conditions for almost sure GUAS and GES. An illustrative example will be proposed to demonstrate the reasonableness of MUSFs and its effectiveness in the stability analysis of randomly switching time-varying systems.
@@ III-A An illustrative example
To verify stability, we introduce the Lyapunov exponent [9]:
$\lim_{t\rightarrow\infty}\frac{1}{t}\ln|x(t)|$
@@
Let’s consider that all the functions $a_{i}(s)$ are constants, that is $a_{i}(s)=\bar{a}_{i}\in\mathbb{R}$ for all $i\in\mathbb{S}$, then $\int_{t_{i_{k}}}^{t_{i_{k+1}}}a_{i}(s)\mathrm{d}s=\bar{a}_{i}(t_{i_{k+1}}-t_{i_{k}})=\bar{a}_{i}S_{i}(k)$. By the strong law of large number [34] and ergodicity of $r$, we have
$\lim_{t\rightarrow\infty}\sum_{i\in\mathbb{S}}\frac{N_{i}(t,0)}{t}\frac{1}{N_{i}(t,0)}\sum_{k=0}^{N_{i}(t,0)}a_{i}S_{i}(k)$
$=$ $\sum_{i\in\mathbb{S}}\frac{\pi_{i}}{m_{i}}a_{i}m_{i}=\sum_{i\in\mathbb{S}}\pi_{i}a_{i},\ \mathrm{a.s.}$ (11)
If $\sum_{i\in\mathbb{S}}\pi_{i}a_{i}<0$, system will be GES a.s. In the above example, irreducibility of the semi-Markov process does play a critical in deriving almost sure stability. Meanwhile, we can notice that for each mode $i$ sequence $\left\{\bar{a}_{i}S_{i}(k)\right\}_{k\in\mathbb{N}^{+}}$ is an independent, identically distributed (i.i.d.) process and also a wide-sense stationary process. This fact also plays a crucial role in deducing (III-A), as we will discuss next.
For each $i$, we define a random variable for sequence $\{\bar{a}_{i}S_{i}(k)\}$ which is its time average:
@@
We then use Chebychev’s inequality to verify (III-A) holds:
@@
$\mathrm{Var}(A_{S}^{i})$
$=$ $\mathbb{E}\left[(A_{S}^{i}-\bar{A}_{S}^{i})^{2}\right]$
@@
$\lim_{t\rightarrow\infty}\frac{1}{t}\ln|x(t)|$
$=$ $\lim_{t\rightarrow\infty}\sum_{i\in\mathbb{S}}\frac{N_{i}(t,0)}{t}\frac{1}{N_{i}(t,0)}\sum_{k=0}^{N_{i}(t,0)}\int_{t_{i_{k}}}^{t_{i_{k+1}}}a_{i}(s)\mathrm{d}s,$
@@
Same as proving (III-A), we can define a random variable $A_{y}^{i}$ from $\{Y_{k}^{i}\}_{k\in\mathbb{N}^{+}}$ as $A_{y}^{i}=\lim_{N_{i}\rightarrow\infty}\frac{1}{N_{i}}\sum_{k=0}^{N_{i}}Y_{k}^{i}$ and $\bar{A}_{y}^{i}=\mathbb{E}[Y_{k}^{i}]=M_{i}$, where $N_{i}=N_{i}(t,t_{0})$, $N_{i}\rightarrow\infty$ as $t\rightarrow\infty$ and prove that the sequence $\{Y^{i}_{k}\}_{k\in\mathbb{N}^{+}}$ is ergodic by the Chebychev’s inequality. The Lyapunov exponent (10) follows from the ergodicity of $\{Y^{i}_{k}\}_{k\in\mathbb{N}^{+}}$ that
$\lim_{t\rightarrow\infty}\frac{1}{t}\ln|x(t)|\leq$ $\lim_{t\rightarrow\infty}\sum_{i\in\mathbb{S}}\frac{N_{i}(t,0)}{t}\frac{1}{N_{i}(t,0)}\sum_{k=0}^{N_{i}(t,0)}b_{i}(S_{i}(k))$
@@
When $\sum_{i\in\mathbb{S}}\frac{\pi_{i}}{m_{i}}M_{i}<0$, system is GES a.s. We derive a sufficient conditions of GES a.s. for time-varying systems.
@@ III-B Indefinite multiple Lyapunov functions and mean uniformly functions
The iMLFs scheme is given as follow.
@@
    where $\dot{V}_{i}(t,x)=\frac{\partial V_{i}(t,x)}{\partial t}+\frac{\partial V_{i}(t,x)}{\partial x}f_{i}(t,x)$;
@@
Time-varying scalar function $\lambda(t)\in\mathcal{PC}(\mathbb{R}^{+};\mathbb{R})$ is said to be a mean uniformly stable function (MUSF), if for corresponding real-valued random variable $S>0$ with $\mathbb{E}[S]<\infty$, there exists a continuous function $\varphi(S)\in\mathcal{C}(\mathbb{R}^{+};\mathbb{R})$ such that $\mathbb{E}\left[\varphi(S)\right]<\infty$ and
$\int_{t}^{t+S}\lambda(h)\mathrm{d}h\leq\varphi(S),\ \forall t\geq t_{0}.$ (16)
@@
When applying the MUSF condition within the iMLF framework, we select the sojourn time $S_{i}$ as the corresponding random variable for $\lambda_{i}(t)$. This allows us to compute the integral $\int_{t}^{t+S_{i}}\lambda_{i}(h)\mathrm{d}h$ and derive the corresponding function $\varphi_{i}(\cdot)$. In this process, we treat $S_{i}$ as a coefficient in the definite integral calculation, yielding $\varphi_{i}(S_{i})$. Then, by taking the expectation of both $\int_{t}^{t+S_{i}}\lambda_{i}(h)\mathrm{d}h$ and $\varphi_{i}(S_{i})$, we obtain a real-valued bound that captures the cumulative variations of the time-varying function $\lambda_{i}(t)$ over the random sojourn time $S_{i}$. Due to the piecewise continuity of function $\lambda_{i}(t)$, its integral is continuous, and existence of a continuous function $\varphi_{i}(\cdot)$ bounding this integral is reasonable. Meanwhile, function $\varphi_{i}$ is independent on time but dependent on sojourn time intervals, which performs the uniformity in time. $\mathbb{E}[\int_{t}^{t+S_{i}}\lambda_{i}(h)\mathrm{d}h]\leq\mathbb{E}[\varphi_{i}(S_{i})]$ implies that all subsystems are mean bounded during their respective sojourn time intervals. Thus, we name this constraint for time-varying scalar function $\lambda_{i}(t)$ as ‘mean uniformly stable function’.